# What is computable and non-computable in the quantum domain:

### v1

@@ Abstract
Recent progress in developing computational devices based on quantum effects and demonstrations of solving various tasks using them has actualized the question of the origin of the quantum advantage. Although various attempts to quantify and characterize the nature of quantum computational advantage have been made, this question in the general context remains open: There is no universal approach that helps to define a scope of problems that quantum computers are able to speed up, theoretically and in practice. Here we review an approach to this question based on the concept of complexity and reachability of quantum states. On the one hand, the class of quantum states that is of interest for quantum computing should be complex, i.e. non-simulatable with classical computers with less than exponential resources. On the other hand, such quantum states should be reachable on a practical quantum computer. This means that a unitary corresponding to the transformation of quantum states from initial to desired can be decomposed in a sequence of single- and two-qubit gates with of no more than polynomial in the number of qubits. Our consideration paves the way towards understanding the scope of problems that can be solved by a quantum computer by formulating a sequence of statements and conjectures on various sets of quantum states.
@@ I Introduction
Historically, the Feynman’s approach to building quantum computers is to use such devices to simulate other quantum systems Feynman (1982, 1986) that are believed to be hard to simulate using classical resources. Specifically, it is conjectured that for a class of quantum many-body systems resources to simulate it, i.e., predict the probability of a given measurement outcome (a so-called strong simulation) or mimic a sampling measurement outcomes (weak simulation), grows exponentially with the system size. This is known as the entanglement frontier Preskill (2012). For example, let us consider a state of $n$ qubits, where each of the qubit is described by the state |$\ket{\psi}=\bigotimes_{i}(\alpha_{i}\ket{0}+\beta_{i}\ket{1})$, where $\alpha_{i}$ and $\beta_{i}$ are complex numbers with the normalization condition $|\alpha_{i}|^{2}+|\beta_{i}|^{2}=1$ for $i=1,\ldots,n$. The description of such a state requires $2n$ real numbers, if, e.g., Bloch sphere angles are used. However, if we apply a long enough sequence of single-qubit operations and two-qubit entangling operations, e.g., the ${\rm CNOT}$ gate Ladd et al. (2010), between different pairs of qubits, then its many-body quantum states becomes entangled, and there will no obvious way to simulate such a state with linear resources. One may conjecture that we would need resources scaling as $2^{n}$, i.e. exponentially with the system size. For the $n=100$ case, a straightforward simulation requires storing $2^{100}$-dimensional complex vector in the memory and calculating the results of rotations in $2^{100}$ size space, which seems to be impossible with the use of any foreseen classical computing device. From this perspective, having any entangled superposition state can be considered as the condition for achieving quantum computational advantage.
However, the celebrated Gottesman-Knill theorem Gottesman (1998); Nielsen and Chuang (2000); Aaronson and Gottesman (2004) says that sometimes it is not the case. A many-body entangled state that is prepared by set of gate belonging to the Clifford set applied to a computational basis state, the so-called stabilizer state, is amenable to polynomial-resource strong simulation with respect to any Pauli measurement, including the computational basis measurement. An example of entangled many-body quantum states that belongs to this class are the Greenberger-Horne-Zeilinger (GHZ) state and graph states. This consideration shows that a way of thinking about quantum computers as devices, whose computational advantage is related only to the entangled superposition nature of the underlying quantum state, is naive. Another example is quantum circuits consisting matchgates, which is known to be efficiently simulatable on a classical computer Ermakov et al. (2024).
@@ II Quantum state picture and universal gate-based quantum computing model
The main object of interest in our case is an $n$-qubit quantum state |$\ket{\Psi_{n}}$, which appears after execution of all the unitaries applied to |$\ket{0}^{\otimes n}$ in a quantum circuit, i.e. the quantum state appearing before computational basis measurements. From different prospectives one may think not about complex quantum states, but complex quantum processes, complex Hamiltonians that govern non-trivial quantum dynamics, or complex measurements. In the first case, the Jamiolkowski Isomorphism allows considering all quantum processes, i.e. ones correspond to quantum circuits, as quantum states in the extended Hilbert space. In the case of complex Hamiltonians and measurements, the equivalence between adiabatic Aharonov et al. (2007), measurement-based Raussendorf and Briegel (2001), and gate-based Ladd et al. (2010) quantum computing models allows one to consider the complexity associated specifically with quantum states without loss of generality. We note that an arbitrary readout measurement can be always reduced to the computational basis measurement by adding the corresponding unitary to the circuit. A similar trick can be applied to the consideration of measurement-based feed-forward operations, which can be replaced by controlled unitary gates and postponed measurements. Thus, our approach based on the consideration of the complexity of quantum states is universal in the view of universality of the gate-based model of quantum computing.

### v2

@@ Abstract
Significant advances in the development of computing devices based on quantum effects and the demonstration of their use to solve various problems have rekindled interest in the nature of the “quantum computational advantage”. Although various attempts to quantify and characterize the nature of the quantum computational advantage have previously been made, this question largely remains open. Indeed, there is no universal approach that allows determining the scope of problems whose solution can be accelerated by quantum computers, in theory of in practice. In this paper, we consider an approach to this question based on the concept of complexity and reachability of quantum states. On the one hand, the class of quantum states that are of interest for quantum computing must be complex, i.e., not amenable to simulation by classical computers with less than exponential resources. On the other hand, such quantum states must be reachable on a practically feasible quantum computer. This means that the unitary operation that transforms the initial quantum state into the desired one must be decomposable into a sequence of one- and two-qubit gates of a length that is at most polynomial in the number of qubits. By formulating several statements and conjectures, we discuss the question of describing a class of problems whose solution can be accelerated by a quantum computer.
@@ I Introduction
Historically, the Feynman’s approach to building quantum computers is to use such devices to simulate other quantum systems Feynman (1982, 1986) that are believed to be hard to simulate using classical resources. Specifically, it is conjectured that for a class of quantum many-body systems resources to simulate it, i.e., predict the probability of a given measurement outcome (a so-called strong simulation) or mimic a sampling measurement outcomes (weak simulation), grows exponentially with the system size. This is known as the entanglement frontier Preskill (2012). For example, let us consider a state of $n$ qubits, where each of the qubit is described by the state |$\ket{\psi}=\bigotimes_{i}(\alpha_{i}\ket{0}+\beta_{i}\ket{1})$, where $\alpha_{i}$ and $\beta_{i}$ are complex numbers with the normalization condition $|\alpha_{i}|^{2}+|\beta_{i}|^{2}=1$ for $i=1,\ldots,n$. The description of such a state requires up to $2n$ real numbers, if, e.g., Bloch sphere angles are used for parametrization. However, if we apply a long enough sequence of single-qubit operations and two-qubit entangling operations, e.g., the ${\rm CNOT}$ gate Ladd et al. (2010), between different pairs of qubits, then its many-body quantum state becomes entangled. It is then seems that there will no obvious way to simulate such a state with linear resources. One may conjecture that we would need resources scaling as $2^{n}$, i.e. exponentially with the system size. For the $n=100$ case, a straightforward simulation requires storing $2^{100}$-dimensional complex vector in the memory and calculating the results of rotations in $2^{100}$ size space, which seems to be impossible with the use of any foreseen classical computing device. From this perspective, having any entangled superposition state can be considered as the condition for achieving quantum computational advantage.
However, the celebrated Gottesman-Knill theorem Gottesman (1998); Nielsen and Chuang (2000); Aaronson and Gottesman (2004) says that sometimes it is not the case. A many-body entangled state that is prepared by set of gates belonging to the Clifford set applied to a computational basis state, the so-called stabilizer state, is amenable to polynomial-resource strong simulation with respect to any Pauli measurement, including the computational basis measurement. An example of entangled many-body quantum states that belongs to this class are the Greenberger-Horne-Zeilinger (GHZ) state and graph states. This consideration shows that a way of thinking about quantum computers as devices, whose computational advantage is related only to the entangled superposition nature of the underlying quantum state, is naive. Another example is quantum circuits consisting matchgates, which is known to be efficiently simulatable on a classical computer Ermakov et al. (2024).
@@ II Quantum state picture and universal gate-based quantum computing model
The main object of interest in our case is an $n$-qubit quantum state |$\ket{\Psi_{n}}$, which appears after execution of all the unitaries applied to |$\ket{0}^{\otimes n}$ in a quantum circuit, i.e. the quantum state appearing before computational basis measurements. From a different prospective, one may think not about complex quantum states, but complex quantum processes, complex Hamiltonians that govern non-trivial quantum dynamics, or complex measurements. In the first case, the Jamiolkowski Isomorphism allows considering all quantum processes, i.e. ones correspond to quantum circuits, as quantum states in the extended Hilbert space. In the case of complex Hamiltonians and measurements, the equivalence between adiabatic Aharonov et al. (2007), measurement-based Raussendorf and Briegel (2001), and gate-based Ladd et al. (2010) quantum computing models allows one to consider the complexity associated specifically with quantum states without loss of generality. We note that an arbitrary readout measurement can be always reduced to the computational basis measurement by adding the corresponding unitary to the circuit. A similar trick can be applied to the consideration of measurement-based feed-forward operations, which can be replaced by controlled unitary gates and postponed measurements. Thus, our approach based on the consideration of the complexity of quantum states is universal in the view of universality of the gate-based model of quantum computing.